Perpetual Inventory Using FIFO Beginning inventory, purchases, and sales for Item PK95 are as follows: July 1 Inventory 55 units @ $23 9 Sale 43 units 13 Purchase 49 units @ $26 28 Sale 17 units Assuming a perpetual inventory system and using the first-in, first-out (FIFO) method, determine (a) the cost of merchandise sold on July 28 and (b) the inventory on July 31. a. Cost of merchandise sold on July 28 $ b. Inventory on July 31 $
Answer #1
a
Merchandise units sold on July 28:
From July 1 Inventory 12 =55-43
From July 13 Purchase 5 =17-12
Cost of merchandise sold on July 28 406 =(12*23)+(5*26)
b
Inventory on July 31 1144 =(49-5)*26